Jen is 26 years younger than her mom. Together their age is 0. hOW OLD IS jEN'S MOM?
Answer by stanbon(75887)

Jen is 26 years younger than her mom. Together their age is 0.
HOW OLD IS jEN'S MOM?
----
Equations:
j = m - 26
j + m = 0
--------------
Substitute for "j" and solve for "m"::
m-26 + m = 0
2m = 26
m = 13 yrs
-----
The conditions are contradictory.
Cheers,
Stan H.